# Robust methods for Darcy–Forchheimer equationsDas, Hutridurga, Pani, Ruiz-Baier

### v1

@@ 3 Continuous robust solvability
$\|p\|_{\mathrm{Q}}$ $\leq\frac{1}{\beta}\|B^{*}(p)\|_{\mathbf{V}}\leq\|{F}\|_{\mathbf{V}^{\prime}}+\|\{A+C\}(\boldsymbol{u})\|_{\mathbf{V}^{\prime}}.$
Then, we employ again the local Lipschitz continuity of $A+C$ setting $\boldsymbol{v}=\boldsymbol{0}$ in (3.7), and appeal to the inequality (3.11). Thus we arrive at
$\|p\|_{\mathrm{Q}}$ $\leq\|{F}\|_{\mathbf{V}^{\prime}}+\|\boldsymbol{u}\|_{\mathbf{V}}+\|\boldsymbol{u}\|_{\mathbf{V}}^{2}$
@@
$\leq\|{F}\|_{\mathbf{V}^{\prime}}+\frac{1}{\alpha}\bigl\{\|{F}\|_{\mathbf{V}^{\prime}}+\frac{1+\alpha}{\beta}\|G\|_{\mathrm{Q}^{\prime}}+\frac{1}{\beta^{2}}\|G\|_{\mathrm{Q}^{\prime}}^{2}\bigr\}$
$\quad+\frac{1}{\alpha^{2}}\bigl\{\|{F}\|_{\mathbf{V}^{\prime}}+\frac{1+\alpha}{\beta}\|G\|_{\mathrm{Q}^{\prime}}+\frac{1}{\beta^{2}}\|G\|_{\mathrm{Q}^{\prime}}^{2}\bigr\}^{2}$
@@
$=\frac{1}{{\alpha}}\bigl\{(1+\alpha)\|{F}\|_{\mathbf{V}^{\prime}}+\frac{1+\alpha}{\beta}\|G\|_{\mathrm{Q}^{\prime}}+\frac{1}{\beta^{2}}\|G\|_{\mathrm{Q}^{\prime}}^{2}\bigr\}$
@@
$\quad+\frac{1}{\alpha^{2}}\bigl\{\|{F}\|_{\mathbf{V}^{\prime}}+\frac{1+\alpha}{\beta}\|G\|_{\mathrm{Q}^{\prime}}+\frac{1}{\beta^{2}}\|G\|_{\mathrm{Q}^{\prime}}^{2}\bigr\}^{2}.$
This concludes the rest of the proof.
@@ 4 Mixed finite element method and robust solvability
In addition, let $\mathbb{RT}_{k}(\mathcal{T}_{h})$} denote the $\mathbf{H}_{\Gamma_{\boldsymbol{u}}}({\mathop{\mathrm{div}}\nolimits},\Omega)$-discrete conforming subspace defined by the Raviart–Thomas elements of order $k$, i.e.,
@@
where we denote by $\widetilde{\mathbb{P}}_{k}(K)$ the space of all polynomials over $K$ of degree exactly equal to $k$ and $^{\prime}\oplus^{\prime}$ denotes the direct sum of spaces. With this we define the discrete weighted space for the velocity as follows
(4.1) $\mathbf{V}\supseteq\mathbf{V}_{h}:=\kappa^{-\frac{1}{2}}\mathbf{L}_{h}^{2}(\Omega)\cap\mathbf{H}_{h,\Gamma_{\boldsymbol{u}}}(\mathop{\mathrm{div}}\nolimits,\Omega)\cap\mathtt{F}^{\frac{1}{3}}\mathbf{L}_{h}^{3}(\Omega),$
where ${\mathbf{L}}_{h}^{2}(\Omega),{\mathbf{H}}_{{h,\Gamma_{\boldsymbol{u}}}}({\mathop{\mathrm{div}}\nolimits},\Omega),$ and ${\mathbf{L}}_{h}^{3}(\Omega)$ correspond to functions in $\mathbb{RT}_{k}(\mathcal{T}_{h})$} equipped with the ${\mathbf{L}}^{2}(\Omega),{\mathbf{H}}({\mathop{\mathrm{div}}\nolimits},\Omega),$ and $\mathbf{L}^{3}(\Omega)$-norms, respectively.
In turn, define the discrete gradient operator $\nabla_{h}:{\mathbb{P}_{k}(\mathcal{T}_{h})}\to{\mathbb{RT}_{k}(\mathcal{T}_{h})}$ by (c.f., [2, eqn. (3.1)])
@@
${\mathrm{W}^{1,\frac{3}{2}}}(\mathcal{T}_{h})$ $:=\{p\in\mathrm{L}^{2}(\Omega):p|_{K}\in\mathrm{W}^{1,\frac{3}{2}}(K),\quad\forall K\in\mathcal{T}_{h}\},$
and whenever needed we add a subscript $\Gamma_{i}$ denoting the part on the boundary where the value of the function is prescribed to zero. These spaces are equipped with broken Sobolev seminorms $|\bullet|_{{1,\mathcal{T}_{h}}}$ and $|\bullet|_{{1,\frac{3}{2},\mathcal{T}_{h}}}$, defined as
${\mathrm{H}^{1}}(\mathcal{T}_{h})\ni\zeta$ $\mapsto|\zeta|_{{1,\mathcal{T}_{h}}}:=\biggl(\sum_{K\in\mathcal{T}_{h}}\int_{K}|{\nabla}\zeta|^{{2}}\biggr)^{\frac{1}{2}},$
$\mathrm{W}^{1,\frac{3}{2}}(\mathcal{T}_{h})\ni\xi$ $\mapsto|\xi|_{{1,\frac{3}{2},\mathcal{T}_{h}}}:=\biggl(\sum_{K\in\mathcal{T}_{h}}\int_{K}|{\nabla}\xi|^{\frac{3}{2}}\biggr)^{\frac{2}{3}},$
respectively.
Furthermore, we define the following weighted space for pressure, which is a sum of the infinite-dimensional (but broken) spaces $\mathrm{L}^{2}(\Omega)$, $\mathrm{H}^{1}_{\Gamma_{p}}(\mathcal{T}_{h})$ and $\mathrm{W}^{1,\frac{3}{2}}_{\Gamma_{p}}(\mathcal{T}_{h})$:
(4.3) ${\widehat{\mathrm{Q}}}(\mathcal{T}_{h}):=\mathrm{L}^{2}(\Omega)+\kappa^{\frac{1}{2}}\mathrm{H}^{1}_{\Gamma_{p}}(\mathcal{T}_{h})+\mathtt{F}^{-\frac{1}{3}}\mathrm{W}^{1,\frac{3}{2}}_{\Gamma_{p}}(\mathcal{T}_{h}).$
@@
$\|q\|_{{\widehat{\mathrm{Q}}(\mathcal{T}_{h})}}:=\inf_{\begin{subarray}{c}q=q_{1}+q_{2}+q_{3}\\ (q_{1},q_{2},q_{3})\in\\ \mathrm{L}^{2}(\Omega)\times\mathrm{H}_{\Gamma_{p}}^{1}(\mathcal{T}_{h})\times\mathrm{W}_{\Gamma_{p}}^{1,\frac{3}{2}}(\mathcal{T}_{h})\end{subarray}}\left(\|q_{1}\|_{0,\Omega}+\kappa^{\frac{1}{2}}\,|q_{2}|_{1,\mathcal{T}_{h}}+\mathtt{F}^{-\frac{1}{3}}|q_{3}|_{1,\frac{3}{2},\mathcal{T}_{h}}\right).$
Finally, we efine a discrete pressure space as
(4.4) $\mathrm{Q}_{h}:=\kappa^{\frac{1}{2}}\mathrm{H}^{1}_{h,{\Gamma_{p}}}(\Omega)+\mathrm{L}_{h}^{2}(\Omega)+\mathtt{F}^{-\frac{1}{3}}\mathrm{W}^{1,\frac{3}{2}}_{h,\Gamma_{p}}(\Omega),$
where $\mathrm{L}_{h}^{2}(\Omega),\mathrm{H}^{1}_{h}(\Omega),$ and $\mathrm{W}^{1,\frac{3}{2}}_{h}(\Omega)$ are polynomial spaces having discrete functions from $\mathbb{P}_{k}(\mathcal{T}_{h})$}, and equipped with the $\mathrm{L}^{2}(\Omega)$ norm and the broken $\mathrm{H}^{1}(\mathcal{T}_{h}),\,\mathrm{W}^{1,\frac{3}{2}}(\mathcal{T}_{h})$ seminorms, respectively.
@@
We equip $\mathrm{Q}_{h}$ with the $\|\bullet\|_{\widehat{\mathrm{Q}}(\mathcal{T}_{h})}$ norm, and observe that $\mathrm{Q}_{h}\subseteq\widehat{\mathrm{Q}}(\mathcal{T}_{h})$, but $\mathrm{Q}_{h}$ is not a subspace of $\mathrm{Q}$. This entails a non-conforming (in the pressure) scheme, requiring to define a discrete bilinear form associated with the discrete divergence operator $b_{h}:\mathbf{V}_{h}\times\mathrm{Q}_{h}\to\mathbb{R}$ as well as a discrete linear functional $G_{h}:{\mathrm{Q}_{h}}\to\mathbb{R}$, as follows
$b_{h}(\boldsymbol{v}_{h},q_{h}):=\int_{\Omega}\mathop{\mathrm{div}}\nolimits\boldsymbol{v}_{h}\,q_{h},\quad G_{h}(q_{h}):=\int_{\Omega}g\,q_{h},\quad\forall\boldsymbol{v}_{h}\in\mathbf{V}_{h},q_{h}\in{\mathrm{Q}_{h}}.$
@@
Similarly to the continuous case, an analysis using $\mathbf{V}_{h}=(\mathbb{RT}_{k}(\mathcal{T}_{h}),\|\bullet\|_{\mathbf{H}^{3}(\mathop{\mathrm{div}}\nolimits,\Omega)})$ and $\mathrm{Q}_{h}=(\mathbb{P}_{k}(\mathcal{T}_{h}),\|\bullet\|_{\mathrm{L}^{2}(\Omega)})$ can be performed (using, e.g., [12, Theorem 4.1], [31, Theorem 3.5]) however the coercivity on the kernel and inf-sup conditions (and therefore also the stability estimates) are not necessarily robust in $\kappa,\mathtt{F}$. To achieve such robustness we employ the weighted spaces (4.1) and (4.4).

### v2

@@ 3 Continuous robust solvability
$B({\boldsymbol{u}^{m}})=G\quad\text{ and }\quad\|{\boldsymbol{u}^{m}}\|_{\mathbf{V}}\leq\frac{1}{\beta}\|G\|_{\mathrm{Q}^{\prime}}.$
Now, for the uniformly strong monotonicity of $A+C$ (cf. (3.8)), we choose $\boldsymbol{u}=\widetilde{\boldsymbol{u}},\boldsymbol{v}=\boldsymbol{0}$ and $\boldsymbol{w}=-{\boldsymbol{u}^{m}}$, and then for equation (2.3a), we set $\boldsymbol{v}=\widetilde{\boldsymbol{u}}$. Further, we use the local Lipschitz continuity of $A$, by setting $\boldsymbol{v}=\boldsymbol{0}$ in (3.7), to conclude that
$\alpha\|\widetilde{\boldsymbol{u}}\|^{2}_{\mathbf{V}}$ $\leq\langle\{A+C\}(\widetilde{\boldsymbol{u}}+{\boldsymbol{u}^{m}})-\{A+C\}(\boldsymbol{0}+{\boldsymbol{u}^{m}}),\widetilde{\boldsymbol{u}}\rangle$
$\leq(\|{F}\|_{\mathbf{V}^{\prime}}+\|\{A+C\}({\boldsymbol{u}^{m}})\|_{\mathbf{V}^{\prime}})\|\widetilde{\boldsymbol{u}}\|_{\mathbf{V}}$
$\leq(\|{F}\|_{\mathbf{V}^{\prime}}+\|{\boldsymbol{u}^{m}}\|_{\mathbf{V}}+\|{\boldsymbol{u}^{m}}\|_{\mathbf{V}}^{r-1})\|\widetilde{\boldsymbol{u}}\|_{\mathbf{V}}$
@@
$\leq\bigl(\|{F}\|_{\mathbf{V}^{\prime}}+\frac{1}{\beta}\|G\|_{\mathrm{Q}^{\prime}}+\frac{1}{\beta^{r-1}}\|G\|_{\mathrm{Q}^{\prime}}^{r-1}\bigr)\|\widetilde{\boldsymbol{u}}\|_{\mathbf{V}},$
and hence, using also the triangle inequality, we find that
@@
(3.10) $\|\boldsymbol{u}\|_{\mathbf{V}}\leq\|\widetilde{\boldsymbol{u}}\|_{\mathbf{V}}+\|{\boldsymbol{u}^{m}}\|_{\mathbf{V}}\leq\frac{1}{\alpha}\bigl(\|{F}\|_{\mathbf{V}^{\prime}}+\frac{1+\alpha}{\beta}\|G\|_{\mathrm{Q}^{\prime}}+\frac{1}{\beta^{r-1}}\|G\|_{\mathrm{Q}^{\prime}}^{r-1}\bigr).$
On the other hand, from (3.9) and employing again the local Lipschitz continuity of $A+C$ with $\boldsymbol{v}=\boldsymbol{0}$ in (3.7), we can invoke inequality (3.10) to arrive at
$\|p\|_{\mathrm{Q}}$ $\leq\|{F}\|_{\mathbf{V}^{\prime}}+\|\boldsymbol{u}\|_{\mathbf{V}}+\|\boldsymbol{u}\|_{\mathbf{V}}^{r-1}$
@@
$\leq\|{F}\|_{\mathbf{V}^{\prime}}+\frac{1}{\alpha}\bigl(\|{F}\|_{\mathbf{V}^{\prime}}+\frac{1+\alpha}{\beta}\|G\|_{\mathrm{Q}^{\prime}}+\frac{1}{\beta^{r-1}}\|G\|_{\mathrm{Q}^{\prime}}^{r-1}\bigr)$
$\quad+\frac{1}{\alpha^{2}}\bigl(\|{F}\|_{\mathbf{V}^{\prime}}+\frac{1+\alpha}{\beta}\|G\|_{\mathrm{Q}^{\prime}}+\frac{1}{\beta^{r-1}}\|G\|_{\mathrm{Q}^{\prime}}^{r-1}\bigr)^{r-1}$
$=\frac{1}{{\alpha}}\bigl((1+\alpha)\|{F}\|_{\mathbf{V}^{\prime}}+\frac{1+\alpha}{\beta}\|G\|_{\mathrm{Q}^{\prime}}+\frac{1}{\beta^{r-1}}\|G\|_{\mathrm{Q}^{\prime}}^{r-1}\bigr)$
$\quad+\frac{1}{\alpha^{2}}\bigl(\|{F}\|_{\mathbf{V}^{\prime}}+\frac{1+\alpha}{\beta}\|G\|_{\mathrm{Q}^{\prime}}+\frac{1}{\beta^{r-1}}\|G\|_{\mathrm{Q}^{\prime}}^{r-1}\bigr)^{r-1}.$
This concludes the rest of the proof. Note that if $|\Gamma_{p}|=0$ then the weighted spaces take the form
$\mathbf{V}=\kappa^{-\frac{1}{2}}\mathbf{L}^{2}(\Omega)\cap\mathbf{H}_{0}(\mathop{\mathrm{div}}\nolimits,\Omega)\cap{\mathtt{F}^{\frac{1}{r}}\mathbf{L}^{r}(\Omega)},\quad\mathrm{Q}=\mathrm{L}_{0}^{2}(\Omega)+\kappa^{\frac{1}{2}}\widetilde{\mathrm{H}}^{1}(\Omega)+\mathtt{F}^{-\frac{1}{r}}\widetilde{\mathrm{W}}^{1,r^{\prime}}(\Omega),$
where $\mathrm{L}_{0}^{2}(\Omega)=\{q\in\mathrm{L}^{2}(\Omega):\int_{\Omega}q=0\}$, $\widetilde{\mathrm{H}}^{1}(\Omega)=\{w\in\mathrm{H}^{1}(\Omega):\int_{\Omega}w=0\}$, and $\widetilde{\mathrm{W}}^{1,r^{\prime}}(\Omega)=\{z\in\mathrm{W}^{1,r^{\prime}}(\Omega):\int_{\Omega}z=0\}$. On the other hand, if $|\Gamma_{\boldsymbol{u}}|=0$ then
$\mathbf{V}=\kappa^{-\frac{1}{2}}\mathbf{L}^{2}(\Omega)\cap\mathbf{H}(\mathop{\mathrm{div}}\nolimits,\Omega)\cap{\mathtt{F}^{\frac{1}{r}}\mathbf{L}^{r}(\Omega)},\quad\mathrm{Q}=\mathrm{L}^{2}(\Omega)+\kappa^{\frac{1}{2}}\mathrm{H}_{0}^{1}(\Omega)+\mathtt{F}^{-\frac{1}{r}}\mathrm{W}_{0}^{1,r^{\prime}}(\Omega).$
@@ 4 Mixed finite element method and robust solvability
In addition, let $\mathbb{RT}_{k}(\mathcal{T}_{h})$} denote the discrete $\mathbf{H}_{\Gamma_{\boldsymbol{u}}}({\mathop{\mathrm{div}}\nolimits},\Omega)$-conforming subspace defined by the Raviart–Thomas elements of order $k$, i.e.,
@@
where we denote by $\widetilde{\mathbb{P}}_{k}(K)$ the space of all polynomials over $K$ of degree exactly equal to $k$ and $^{\prime}\oplus^{\prime}$ denotes the direct sum of spaces. With this we define the discrete weighted space for the velocity as all the vector fields in $\mathbb{RT}_{k}(\mathcal{T}_{h})$ such that their $\mathbf{V}$ norm (cf. (3.2a)) is bounded. Equivalently, and with the aim to emphasise the parameter-dependence, we write this space as
(4.1) $\mathbf{V}\supseteq\mathbf{V}_{h}:=\kappa^{-\frac{1}{2}}\mathbf{L}_{h}^{2}(\Omega)\cap\mathbf{H}_{h,\Gamma_{\boldsymbol{u}}}(\mathop{\mathrm{div}}\nolimits,\Omega)\cap{\mathtt{F}^{\frac{1}{r}}\mathbf{L}_{h}^{r}(\Omega)},$
where ${\mathbf{L}}_{h}^{2}(\Omega),{\mathbf{H}}_{{h,\Gamma_{\boldsymbol{u}}}}({\mathop{\mathrm{div}}\nolimits},\Omega),$ and ${\mathbf{L}}_{h}^{r}(\Omega)$ correspond to functions in $\mathbb{RT}_{k}(\mathcal{T}_{h})$} equipped with the ${\mathbf{L}}^{2}(\Omega),{\mathbf{H}}({\mathop{\mathrm{div}}\nolimits},\Omega),$ and $\mathbf{L}^{r}(\Omega)$-norms, respectively.
In turn, define the discrete gradient operator $\nabla_{h}:{\mathbb{P}_{k}(\mathcal{T}_{h})}\to{\mathbb{RT}_{k}(\mathcal{T}_{h})}$ by (cf., [2, eqn. (3.1)])
@@
${\mathrm{W}^{1,r^{\prime}}}(\mathcal{T}_{h})$ $:=\{p\in\mathrm{L}^{2}(\Omega):p|_{K}\in\mathrm{W}^{1,r^{\prime}}(K),\quad\forall K\in\mathcal{T}_{h}\},$
and whenever needed we add a subscript $\Gamma_{i}$ denoting the part on the boundary where the value of the function is prescribed to zero. These spaces are equipped with DG (broken Sobolev) seminorms $|\bullet|_{{1,\mathcal{T}_{h}}}$ and $|\bullet|_{{1,r^{\prime},\mathcal{T}_{h}}}$, defined as
${\mathrm{H}^{1}}(\mathcal{T}_{h})\ni\zeta$ $\mapsto|\zeta|^{2}_{{1,\mathcal{T}_{h}}}:=\biggl(\sum_{K\in\mathcal{T}_{h}}\|{\nabla}\zeta\|^{2}_{0,K}+\sum_{F\in\mathcal{F}_{h}}h_{F}^{-1}\|[\![\zeta]\!]\|^{2}_{0,F}\biggr)^{\frac{1}{2}},$
$\mathrm{W}^{1,r^{\prime}}(\mathcal{T}_{h})\ni\xi$ $\mapsto{|\xi|_{{1,r^{\prime},\mathcal{T}_{h}}}:=\biggl(\sum_{K\in\mathcal{T}_{h}}\|{\nabla}\xi\|^{r^{\prime}}_{0,r^{\prime},K}+\sum_{F\in\mathcal{F}_{h}}h_{F}^{1-r^{\prime}}\|[\![\xi]\!]\|^{r^{\prime}}_{0,r^{\prime},F}\biggr)^{\frac{1}{r^{\prime}}}},$
respectively, where $\mathcal{F}_{h}$ denotes the set of facets in the mesh $\mathcal{T}_{h}$.
Furthermore, we define the following weighted space for pressure, which is a sum of the infinite-dimensional (but broken) spaces $\mathrm{L}^{2}(\Omega)$, $\mathrm{H}^{1}_{\Gamma_{p}}(\mathcal{T}_{h})$ and $\mathrm{W}^{1,r^{\prime}}_{\Gamma_{p}}(\mathcal{T}_{h})$:
(4.3) ${\widehat{\mathrm{Q}}}(\mathcal{T}_{h}):=\mathrm{L}^{2}(\Omega)+\kappa^{\frac{1}{2}}\mathrm{H}^{1}_{\Gamma_{p}}(\mathcal{T}_{h})+{\mathtt{F}^{-\frac{1}{r}}\mathrm{W}^{1,r^{\prime}}_{\Gamma_{p}}(\mathcal{T}_{h})}.$
@@
(4.4) $\|q\|_{{\widehat{\mathrm{Q}}(\mathcal{T}_{h})}}:=\inf_{\begin{subarray}{c}q=q_{1}+q_{2}+q_{3}\\ (q_{1},q_{2},q_{3})\in\\ \mathrm{L}^{2}(\Omega)\times\mathrm{H}_{\Gamma_{p}}^{1}(\mathcal{T}_{h})\times{\mathrm{W}_{\Gamma_{p}}^{1,r^{\prime}}(\mathcal{T}_{h})}\end{subarray}}\left(\|q_{1}\|_{0,\Omega}+\kappa^{\frac{1}{2}}\,|q_{2}|_{1,\mathcal{T}_{h}}+{\mathtt{F}^{-\frac{1}{r}}|q_{3}|_{1,r^{\prime},\mathcal{T}_{h}}}\right).$
Finally, we define the discrete pressure space as all the functions in $\mathbb{P}_{k}(\mathcal{T}_{h})$ whose $\widehat{\mathrm{Q}}_{h}(\mathcal{T}_{h})$ norm is bounded (see Remark 4.1)
(4.5) $\mathrm{Q}_{h}:=\kappa^{\frac{1}{2}}\mathrm{H}^{1}_{h,{\Gamma_{p}}}(\Omega)+\mathrm{L}_{h}^{2}(\Omega)+{\mathtt{F}^{-\frac{1}{r}}\mathrm{W}^{1,r^{\prime}}_{h,\Gamma_{p}}(\Omega)},$
where $\mathrm{L}_{h}^{2}(\Omega),\mathrm{H}^{1}_{h}(\Omega),$ and $\mathrm{W}^{1,r^{\prime}}_{h}(\Omega)$ are polynomial spaces having discrete functions from $\mathbb{P}_{k}(\mathcal{T}_{h})$}, and equipped with the $\mathrm{L}^{2}(\Omega)$ norm and the broken $\mathrm{H}^{1}(\mathcal{T}_{h}),\,{\mathrm{W}^{1,r^{\prime}}(\mathcal{T}_{h})}$ seminorms, respectively.
@@
We equip $\mathrm{Q}_{h}$ with the $\|\bullet\|_{\widehat{\mathrm{Q}}(\mathcal{T}_{h})}$ norm since we cannot compute, for example, the continuous $\mathrm{H}^{1}(\Omega)$ and $\mathrm{W}^{1,r^{\prime}}(\Omega)$ norms of functions in $\mathbb{P}_{k}(\mathcal{T}_{h})$. We do have that $\mathrm{Q}_{h}\subseteq\widehat{\mathrm{Q}}(\mathcal{T}_{h})$, but $\mathrm{Q}_{h}$ is not a subspace of $\mathrm{Q}$. This entails a non-conforming (in the pressure) scheme, requiring to define a discrete bilinear form associated with the discrete divergence operator $b_{h}:\mathbf{V}_{h}\times\mathrm{Q}_{h}\to\mathbb{R}$ as well as a discrete linear functional $G_{h}:{\mathrm{Q}_{h}}\to\mathbb{R}$, as follows
${b_{h}(\boldsymbol{v}_{h},q_{h}):=-\int_{\Omega}\mathop{\mathrm{div}}\nolimits\boldsymbol{v}_{h}\,q_{h},\quad G_{h}(q_{h}):=-\int_{\Omega}g\,q_{h},}\quad\forall\boldsymbol{v}_{h}\in\mathbf{V}_{h},q_{h}\in{\mathrm{Q}_{h}}.$
@@
Similarly to the continuous case, an analysis using $\mathbf{V}_{h}=(\mathbb{RT}_{k}(\mathcal{T}_{h}),\|\bullet\|_{\mathbf{H}^{{r}}(\mathop{\mathrm{div}}\nolimits,\Omega)})$ and $\mathrm{Q}_{h}=(\mathbb{P}_{k}(\mathcal{T}_{h}),\|\bullet\|_{\mathrm{L}^{2}(\Omega)})$ can be performed (using, e.g., [13, Theorem 4.1], [32, Theorem 3.5]) however the coercivity on the kernel and inf-sup conditions (and therefore also the stability estimates) are not necessarily robust in $\kappa,\mathtt{F}$. To achieve such robustness we employ the weighted spaces (4.1) and (4.5).